Calculate the resultant velocity of an airplane that normally flies at 20 km/h if it encounters a 50-km/h tailwind. If it encounters a 50-km/h headwind
1
Expert's answer
2012-09-28T09:04:37-0400
If it encounters a 50-km/h tailwind, Vr = Va + Vw = 20 + 50 = 70 km/h
If it
encounters a 50-km/h headwind, Vr = Va - Vw = 20 - 50 = - 30 km/h